# To appear in Proceedings of KGML Bridge Program (AAAI-2026) Knowledge-Guided Masked Autoencoder with Linear Spectral Mixing and Spectral-Angle–Aware Reconstruction

### v1

@@ Introduction
Recent advances in self-supervised learning and large-scale transformer architectures have demonstrated impressive representation learning capabilities across vision and multimodal tasks. In particular, Vision Transformer-based Masked Autoencoders (ViT-MAEs) have shown that reconstructing masked input patches provides an effective pretext task for learning transferable features at scale (He et al. 2022; Dosovitskiy et al. 2021). Such architectures have further evolved into large-scale foundation models that can adapt across diverse visual and geospatial applications by leveraging vast unlabeled data (Bommasani et al. 2021). However, purely data-driven reconstruction objectives often overlook domain-specific priors and the physical mechanisms that govern data generation, limiting interpretability and generalization (Willard et al. 2020; Raissi, Perdikaris, and Karniadakis 2019).
@@
In recent years, an emerging paradigm known as knowledge-guided machine learning (KGML) has begun to blur the boundary between purely data-driven methods and science-driven modeling. Rather than treating machine learning as a black box, KGML frameworks strategically integrate domain knowledge (such as physical laws, mixing models, or causally grounded relationships) with representation learning in neural architectures (Karpatne, Jia, and Kumar 2024). For example, Liu et al. (2024) demonstrate how embedding process-based and remote-sensing knowledge into neural frameworks significantly improves soil-carbon estimation across agroecosystems, highlighting the potential of hybrid scientific–data approaches. Drawing on these insights, our work situates itself at the intersection of foundation-model design and domain-informed representation learning.
@@
The endmember matrix $\mathbf{A}$ can be initialized from prior knowledge (e.g., spectral libraries or k-Means prototypes) or learned end-to-end with a small learning rate for adaptive refinement (Ozkan, Kaya, and Akar 2019).
@@
The Spectral Angle Mapper (SAM) offers a complementary perspective to traditional pixel-wise metrics by measuring the angular distance between spectral vectors, independent of their magnitude (Kruse et al. 1993; Yuhas, Goetz, and Boardman 1992). Unlike MSE or Huber Loss, which is sensitive to intensity variations, SAM focuses on preserving the spectral shape, a critical property for identifying materials in hyperspectral data. We define the SAM loss across $N$ samples as:
@@
ID Model
Reconstruction
Loss
Physical
Constraints
Avg.
PSNR ($\uparrow$)
Per-Channel
PSNR (Max) ($\uparrow$)
Avg.
SSIM ($\uparrow$)
Per-Channel
SSIM (Max) ($\uparrow$)
@@
Top-1
Acc(%) ($\uparrow$)
mIoU
(%) ($\uparrow$)
Top-1
Acc(%) ($\uparrow$)
mIoU
(%) ($\uparrow$)
@@
Integrating scientific priors into ML models has been formalized under physics-informed learning and knowledge-guided ML; these approaches provide improved physical realism and extrapolation (Raissi, Perdikaris, and Karniadakis 2019; Willard et al. 2020; Karpatne, Jia, and Kumar 2024; Liu et al. 2024). KGML methods have been used to incorporate support for physical constraints grounded in soil hydrology, including the van Genuchten water retention equations and models (i.e., Richards’ Equation) of hydraulic conductivity (Khandelwal, Pallickara, and Pallickara 2024; Khandelwal et al. 2025); vegetation indices (Bruhwiler et al. 2020); evapotranspiration (Armstrong et al. 2022); preserving graph properties such as betweenness centrality (Matin et al. 2022); masking cloud occlusions in satellite imagery (Khandelwal et al. 2022), accounting for human perceptual limits during visualization (Mitra et al. 2021); masked autoencoders (Faruk et al. 2025; Matin et al. 2025) and accounting for correlations between soil spectroscopic properties (Bachinin et al. 2025; Dey et al. 2025). Our karma framework follows this paradigm by embedding LSMM within a self-supervised transformer framework. Finally, knowledge distillation approaches have been explored in the context of physical phenomena (Matin et al. 2023); our methodology is complementary to these approaches.
@@
LSMM and hyperspectral unmixing have a rich literature; classical formulations and modern deep unmixing approaches show how material signatures can be recovered and used for interpretability (Bioucas-Dias et al. 2012; Ozkan, Kaya, and Akar 2019). karma uses LSMM as an explicit constraint rather than a standalone unmixing network.

### v2

# Knowledge-Guided Masked Autoencoder with Linear Spectral Mixing and Spectral-Angle–Aware Reconstruction

@@ Introduction
Recent advances in self-supervised learning and large-scale transformer architectures have demonstrated impressive representation learning capabilities across vision and multimodal tasks. In particular, Vision Transformer-based Masked Autoencoders (ViT-MAEs) have shown that reconstructing masked input patches provides an effective pretext task for learning transferable features at scale (He et al. 2022; Dosovitskiy et al. 2021). Such architectures have further evolved into large-scale foundation models that can adapt across diverse visual and geospatial applications by leveraging vast unlabeled data (Bommasani et al. 2021). However, purely data-driven reconstruction objectives often overlook domain-specific priors and the physical mechanisms that govern data generation, limiting interpretability and generalization (Willard et al. 2020; Raissi et al. 2019).
@@
In recent years, an emerging paradigm known as knowledge-guided machine learning (KGML) has begun to blur the boundary between purely data-driven methods and science-driven modeling. Rather than treating machine learning as a black box, KGML frameworks strategically integrate domain knowledge (such as physical laws, mixing models, or causally grounded relationships) with representation learning in neural architectures (Karpatne et al. 2024). For example, Liu et al. (2024) demonstrate how embedding process-based and remote-sensing knowledge into neural frameworks significantly improves soil-carbon estimation across agroecosystems, highlighting the potential of hybrid scientific–data approaches. Drawing on these insights, our work situates itself at the intersection of foundation-model design and domain-informed representation learning.
@@
The endmember matrix $\mathbf{A}$ can be initialized from prior knowledge (e.g., spectral libraries or k-Means prototypes) or learned end-to-end with a small learning rate for adaptive refinement (Ozkan et al. 2019).
@@
The Spectral Angle Mapper (SAM) offers a complementary perspective to traditional pixel-wise metrics by measuring the angular distance between spectral vectors, independent of their magnitude (Kruse et al. 1993; Yuhas et al. 1992). Unlike MSE or Huber Loss, which is sensitive to intensity variations, SAM focuses on preserving the spectral shape, a critical property for identifying materials in hyperspectral data. We define the SAM loss across $N$ samples as:
@@
ID Model Reconstruction Loss Physical Constraints Avg. PSNR ($\uparrow$) Per-Channel PSNR (Max) ($\uparrow$) Avg. SSIM ($\uparrow$) Per-Channel SSIM (Max) ($\uparrow$)
@@
Top-1 Acc(%) ($\uparrow$) mIoU (%) ($\uparrow$) Top-1 Acc(%) ($\uparrow$) mIoU (%) ($\uparrow$)
@@
Integrating scientific priors into ML models has been formalized under physics-informed learning and knowledge-guided ML; these approaches provide improved physical realism and extrapolation (Raissi et al. 2019; Willard et al. 2020; Karpatne et al. 2024; Liu et al. 2024). KGML methods have been used to incorporate support for physical constraints grounded in soil hydrology, including the van Genuchten water retention equations and models (i.e., Richards’ Equation) of hydraulic conductivity (Khandelwal et al. 2024, 2025); vegetation indices (Bruhwiler et al. 2020); evapotranspiration (Armstrong et al. 2022); preserving graph properties such as betweenness centrality (Matin et al. 2022); masking cloud occlusions in satellite imagery (Khandelwal et al. 2022), accounting for human perceptual limits during visualization (Mitra et al. 2021); masked autoencoders (Faruk et al. 2025; Matin et al. 2025) and accounting for correlations between soil spectroscopic properties (Bachinin et al. 2025; Dey et al. 2025). Our karma framework follows this paradigm by embedding LSMM within a self-supervised transformer framework. Finally, knowledge distillation approaches have been explored in the context of physical phenomena (Matin et al. 2023); our methodology is complementary to these approaches.
@@
LSMM and hyperspectral unmixing have a rich literature; classical formulations and modern deep unmixing approaches show how material signatures can be recovered and used for interpretability (Bioucas-Dias et al. 2012; Ozkan et al. 2019). karma uses LSMM as an explicit constraint rather than a standalone unmixing network.